# Worldline Formulations of Covariant Fracton Theories

### v1

@@ 5 Conclusion and outlook
To further develop and apply these models, the next natural step involves the construction of the corresponding worldline path integrals. For the tensor model, this is expected to proceed in close analogy with standard worldline models describing spin-$s$ fields, given the similarity in the constraint structure and gauge symmetries. However, it seems not directly feasible to investigate the path integral in configuration space for the vector model (and its deformation), since all classical constraints are quadratic in the momenta and in the bosonic oscillators. This issue requires further investigation. A related issue concerns the possible inclusion of the classical counterpart of the trace constraint in the worldline action. As discussed in the relevant sections, this constraint is essential at the BRST level. However, its inclusion appears to be incompatible with Siegel gauge-fixing. This raises an open question: should the classical action include the trace constraint? Clarifying this point is an important direction for future work.

### v2

@@ 5 Conclusion and outlook
We also aimed to illustrate the connection between worldline and spacetime formulations in a self-consistent way, in order to emphasize the potential of this formulation in a non-trivial example. Indeed, as we discussed, the usual procedure for developing a worldline formulation of a field theory required some extensions, in the present case, in order to take into account the free parameters $\alpha$ and $\beta$, which distinguishes the theories in the family of covariant fracton gauge theories (2.2).
Thus, the result of the paper suggests that the conventional formulation of quantum field theory in terms of worldline in target space extends well beyond the limits in which the latter was initially conceived, and suggests that the worldline formulation may constitute a potential equivalent formulation of quantum field theory.
To further develop and apply the models we studied, the next natural step involves the construction of the corresponding worldline path integrals. For the tensor model, this is expected to proceed in close analogy with standard worldline models describing spin-$s$ fields, given the similarity in the constraint structure and gauge symmetries. However, it seems not directly feasible to investigate the path integral in configuration space for the vector model (and its deformation), since all classical constraints are quadratic in the momenta and in the bosonic oscillators. This issue requires further investigation. A related issue concerns the possible inclusion of the classical counterpart of the trace constraint in the worldline action. As discussed in the relevant sections, this constraint is essential at the BRST level. However, its inclusion appears to be incompatible with Siegel gauge-fixing. This raises an open question: should the classical action include the trace constraint? Clarifying this point is an important direction for future work.
@@
Finally, it has already been observed that the fractonic behaviour can be understood within the “ultralocal” nature of Carrollian physics,222222“Carroll transformations” are the limit of Lorentz transformations by sending the speed of light to zero, [63]. In Carroll limit, the invariant interval $\mathrm{d}s^{2}=-c^{2}\,\mathrm{d}t^{2}+\mathrm{d}x^{2}$, taking the limit as $c$ tends to zero, becomes $\mathrm{d}s^{2}=\mathrm{d}x^{2}$. This means that there exists an absolute space, on which all observers agree. In Carrollian physics, temporal variations dominate over spatial gradients. In the zero-speed limit, light cones are shrunk to the time axis, meaning any causal connection between event points with distinct spatial parts is excluded. In other words, if the speed of light is zero and the only allowed transmissions are those with speed less than that of light, no transmission is allowed, and no event point can be the cause or effect of any event point. For this reason, the Carrollian limit is also called “ultralocal”. since fractons are characterized by reduced mobility due to the conservation of the dipole moment [33]. The possibility of obtaining fractonic Lorentz-breaking models from the covariant formulations has not been explored yet. Now, in [57] the Carrollian contraction of Lorentz-invariant theories is studied in general: it is shown that there always exist two independent contractions of the same Lorentz-invariant theory, which produce two independent Carrollian theories. To do this, one considers the action of the Lorentz-invariant theory in Hamiltonian formulation, in which Lorentz invariance is not manifest. The two limits consist in turning on only the electric or only the magnetic part of the energy, and are distinguished by how the fields are rescaled with respect to the speed of light. We reserve for the future the analysis of the electric/magnetic Carroll contraction of covariant fracton gauge theories.